At a temperature of 3°C, the vapor pressure of a 500g solution of water is 5.7 mmHg. If 150g of of HCl is added, what is the pressure of the solution?
1
Expert's answer
2020-04-24T13:35:08-0400

This task is based on Raoul’s law: "(P_A^0 -P_A)\/P_A^0=X_B" , "P_A^0" - vapor pressure of initial solution, "P_A" - vapor pressure of solution after addition any substance, "X_B" - mole fraction of the subdtance.

"X_B=n_B/\sum n", nB - chemical amount of substance (HCl), n -total amount of all constituents in a solution.

n=m/M , n - chemical amount of substance(mol); m - mass of substance(g); M - molar mass of substance(g/mol). M(H2O)=16G/mol+1*2g/mol=18g/mol; M(HCl)=35,5g/mol+1g/mol=36,5g/mol.

n(H2O)=500g/18g/mol =27,8mol

n(HCl)=150g/36,5g/mol=4,1mol

XB=4,1/(4,1+27,8)=0,13

Express the needed pressure: "P_A=P_A^0(1-X_B)"

PA=5,7mmHg(1-0,13)=4,96mmHg

Answer: 4,96mmHg.